# CLASH: Collaborative Large-Small Hierarchical Framework for Continuous Vision-and-Language Navigation

### v1

@@ VI-C1 Effect of different components affecting the RSMP
Table III presents an ablation of key design choices, revealing several insights: First, leveraging pretrained weights significantly boosts RSMP performance. Both ETPNav (EN) [an2024etpnav] and ScaleVLN (SV) [wang2023scaling] adopt DUET [chen2022think] as their backbone, but differ in data augmentation: EN uses PREVALENT [hao2020towards] (61 MP3D scenes), while SV incorporates an additional 800 scenes from HM3D. As seen in #1 and #5, initializing with SV weights greatly improves generalization on val-unseen. However, combining Habitat training with SV’s large dataset offers diminishing returns and even degrades val-seen performance (#3 and #4), likely due to distribution mismatch. Therefore, we fine-tune only on the standard training set. Second, incorporating depth input improves navigation accuracy (#4 and #6), consistent with EN’s findings. Third, EN removes the local branch from DUET, limiting its ability to exploit pretrained representations. Restoring the dual-branch structure improves cross-modal alignment and boosts performance across both seen and unseen splits (#5 and #6).
@@
# Init.
FT Set
Depth DB R2R-CE Val Seen R2R-CE Val Unseen
@@ VI-C2 Effect of causal learning in RSMP
We integrate the causal learning (CL) modules [wang2024causal, wang2024causality] into the RSMP backbone during VLN-CE fine-tuning to improve generalization. As shown in Table IV, on both R2R-CE and REVERIE-CE, CL improves SPL on both seen and unseen environments. For example, on the REVERIE-CE val-unseen split, SR increases from 30.79 to 35.65 and SPL from 26.41 to 28.46. However, on R2R, the gains are less pronounced, likely because the pretrained SV weights already generalize well and VLN-CE lacks a speaker model [wang2023pasts, wang2023res] to complement instruction diversity. This observation motivates our future work.
@@ VI-C5 Effect of point-goal controller
Table VI compares the rule-based Tryout controller from ETPNav with a fully learned DDPPO-based controller [wijmansdd2019ddppo]. Tryout often leads to deadlocks in cluttered scenes with elevation changes (Fig. 7), whereas DDPPO provides more reliable obstacle avoidance and recovery. Replacing Tryout with DDPPO improves OSR by 2.29% and 3.06%, and SR by 1.25% and 1.97% on the val-seen and val-unseen splits. The slight reduction in SPL is due to DDPPO favoring safer, although sometimes longer, paths instead of the more aggressive shortcuts employed by Tryout.
@@ VI-C7 Analysis on the model efficiency
CMA [krantz_beyond_2020] 37M $2.74\pm 1.32$ $0.03\pm 0.09$ 51 45 34 28
ETPNav [an2024etpnav] 299M $74\pm 0.68 0.0$$07\pm 0.19$ 66 59 57 49
g3d-LF-Mono [wang2025g3d] 763M $2.94\pm 5.08 1.1$$18\pm 0.38$ 47 35 49 37
OpenNav [qiao2024open] 10B $162.41\pm 8.97$ $27.00\pm 3.64$ - - 19 16
NaVID [zhang2024navid] 7B $21.08\pm 15.12$ $0.69\pm 0.92$ 43 38 37 35
UniNaVID [zhang2024uni] 7B $9.93\pm 2.79$ $0.18\pm 0.02$ 58 53 47 43
NaVILA [cheng2024navila] 7B $62.02\pm 65.65$ $0.86\pm 0.17$ 58 53 54 49
StreamVLN [wei2025streamvln] 7B $42.37\pm 41.43$ $0.71\pm 0.16$ 62 58 57 52

### v2

@@ VI-C1 Effect of different components affecting the RSMP
Table III presents an ablation of key design choices, revealing several insights: First, leveraging pretrained weights significantly boosts RSMP performance. Both ETPNav (EN) [3] and ScaleVLN (SV) [69] adopt DUET [11] as their backbone, but differ in data augmentation: EN uses PREVALENT [24] (61 MP3D scenes), while SV incorporates an additional 800 scenes from HM3D. As seen in #1 and #5, initializing with SV weights greatly improves generalization on val-unseen. However, combining Habitat training with SV’s large dataset offers diminishing returns and even degrades val-seen performance (#3 and #4), likely due to distribution mismatch. Therefore, we fine-tune only on the standard training set. Second, incorporating depth input improves navigation accuracy (#4 and #6), consistent with EN’s findings. Third, EN removes the local branch from DUET, limiting its ability to exploit pretrained representations. Restoring the dual-branch structure improves cross-modal alignment and boosts performance across both seen and unseen splits (#5 and #6).
@@
# Init. FT Set Depth DB R2R-CE Val Seen R2R-CE Val Unseen
@@ VI-C2 Effect of causal learning in RSMP
We integrate the causal learning (CL) modules [63, 58] into the RSMP backbone during VLN-CE fine-tuning to improve generalization. As shown in Table IV, on both R2R-CE and REVERIE-CE, CL improves SPL on both seen and unseen environments. For example, on the REVERIE-CE val-unseen split, SR increases from 30.79 to 35.65 and SPL from 26.41 to 28.46. However, on R2R, the gains are less pronounced, likely because the pretrained SV weights already generalize well and VLN-CE lacks a speaker model [61, 57] to complement instruction diversity. This observation motivates our future work.
@@ VI-C5 Effect of point-goal controller
Table VI compares the rule-based Tryout controller from ETPNav with a fully learned DDPPO-based controller [72]. Tryout often leads to deadlocks in cluttered scenes with elevation changes (Fig. 7), whereas DDPPO provides more reliable obstacle avoidance and recovery. Replacing Tryout with DDPPO improves OSR by 2.29% and 3.06%, and SR by 1.25% and 1.97% on the val-seen and val-unseen splits. The slight reduction in SPL is due to DDPPO favoring safer, although sometimes longer, paths instead of the more aggressive shortcuts employed by Tryout.
@@ VI-C7 Analysis on the model efficiency
CMA [38] 37M $2.74\pm 1.32$ $0.03\pm 0.09$ 51 45 34 28
ETPNav [3] 299M $74\pm 0.68 0.0$$07\pm 0.19$ 66 59 57 49
g3d-LF-Mono [66] 763M $2.94\pm 5.08 1.1$$18\pm 0.38$ 47 35 49 37
OpenNav [49] 10B $162.41\pm 8.97$ $27.00\pm 3.64$ - - 19 16
NaVID [77] 7B $21.08\pm 15.12$ $0.69\pm 0.92$ 43 38 37 35
UniNaVID [76] 7B $9.93\pm 2.79$ $0.18\pm 0.02$ 58 53 47 43
NaVILA [13] 7B $62.02\pm 65.65$ $0.86\pm 0.17$ 58 53 54 49
StreamVLN [71] 7B $42.37\pm 41.43$ $0.71\pm 0.16$ 62 58 57 52